OpenStudy (anonymous):

What is the surface area of a great circle of a sphere whose surface area is 44.8 square yeards?

OpenStudy (anonymous):

\[SA(Circle) = 4*\pi*r^2 = 44.8\] \[r = 1.88813949\] Surface Area of the great circle = 2*pi*r \[SA = 11.8635303\]

OpenStudy (anonymous):

Thank you so much!